OpenStudy (tiffanymak1996):

Show that if u(x,y) and v(x,y) each have equal mixed second partials, and if u and v satisfy the Cauchy -Riemann equations, then u, v, and u+v satisfy Laplace's equation.

OpenStudy (tiffanymak1996):

the Cauchy-Riemann eq: du/dx=dv/dy and du/dy=-dv/dx

OpenStudy (tiffanymak1996):

Laplace 's eq: d^2z/dx^2 + d^2z/dy^2=0

OpenStudy (tiffanymak1996):

@zepdrix help?

OpenStudy (tiffanymak1996):

help?

OpenStudy (anonymous):

@tiffanymak1996 what is mean't by that u and v have equal mixed second partials? Is it just implying that u(x,y) and v(x,y) have the same second partial derivative with respect to x or y?

OpenStudy (anonymous):

so ill show you an example for u : using cauchy eq: du/dx = dv/dy take derivative with respect to x : A. d^2u/dx^2 = d^2v/(dxdy) using cauchy second eq: du/dy = -dv/dx take derivative with respect to y: d^2u/dy^2 = -d^2v/(dydx) now v have equal mixed partials so : d^2v/(dydx) = d^2v/(dxdy) so that B. d^2u/dy^2 = -d^2v/(dxdy) using A +B we get: d^2u/dx^2 + d^2u/dy^2 = 0

OpenStudy (tiffanymak1996):

@Coolsector Thanks.